### v4

@@ 9 Conclusion
$\lambda$-eclipse [112] Multiple Subjects PTA Kandinsky v2.2 Contrasstive Learning
MS-Diffusion [56] Multiple Subjects PTA SD Attention-based Operation
ReVersion [115] High-level Semantic TTF SD Regularization
Inv-ReVersion [200] High-level Semantic TTF SD 1.5 Regularization
CusConcept [201] High-level Semantic TTF SD 2.1 Regularization
@@
ADI [117] High-level Semantic TTF SD 2.1 Mask-assisted Generation
PhotoSwap [118] Extra Conditions TTF SD 2.1 Attention-based Operation
PE-VITON [202] Extra Conditions TTF SD Shape and Texture Control
@@
Layout-Control [69] Extra Conditions TTF SD Attention-based Operation
SwapAnything [203] Extra Conditions TTF SD 2.1 Mask-assisted Generation
PE-VITON [202] Extra Conditions TTF SD Shape and Texture Control
@@
Viewpoint Control [122] Extra Conditions TTF SDXL 3D Feature Incorporation
@@
Prompt-Free Diffusion [204] Extra Conditions PTA SD 2.0 Reference Feature Injection
Uni-ControlNet [205] Extra Conditions PTA SD Multi-feature Injection
ViscoNet [206] Extra Conditions PTA SD Multi-feature Injection
@@
Context Diffusion [207] Extra Conditions PTA LDM Multi-feature Injection
FreeControl [208] Extra Conditions PTA SD 1.5; SD 2.1; SDXL Fusion Guidance
@@
Li et al. [209] Extra Conditions PTA SD Attention-based Operation; Mask-assisted Generation
@@
Tune-A-Video [127] Video TTF SD 1.4 Parameter-efficient Fine-tuning
Gen-1 [130] Video TTF LDM Multi-feature Injection
Make-A-Protagonist [210] Video TTF LDM Attention-based Operation
Animate-A-Story [131] Video TTF SD Retrieval-augmented Paradigm
MotionDirector [134] Video TTF ZeroScope; ModelScope Disentanglement Approach
LAMP [132] Video TTF SDXL; SD 1.4 First-frame Conditioned Pipeline
VMC [129] Video TTF Show-1 Parameter-efficient Fine-tuning
SAVE [128] Video TTF VDM Attention-based Operation
Customizing Motion [133] Video TTF ZeroScope Regularization
DreamVideo [135] Video TTF ModelScope Disentanglement Approach
MotionCrafter [211] Video TTF VDM Disentanglement Approach
Customize-A-Video [212] Video TTF ModelScope Multi-stage Refinement
Magic-Me [213] Video TTF SD 1.5; AnimateDiff Multi-stage Refinement
CustomTTT [214] Video TTF VDM Fine-tuning and Alignment
PersonalVideo [215] Video; Face TTF AnimateDiff Identity Injection
@@
CustomVideo [81] Video; Multiple Subjects TTF ZeroScope Data Augmentation; Attention-based Operation
VideoDreamer [80] Video; Multiple Subjects TTF SD 2.1 Data Augmentation
VideoAssembler [125] Video PTA VidRD Reference Feature Injection
VideoBooth [124] Video PTA VDM Reference Feature Injection
VideoMaker [216] Video PTA VDM Reference Feature Injection
SUGAR [217] Video PTA VDM Attention-based Operation
StyleCrafter [54] Video; Style PTA SD 1.5 Reference Feature Injection
ID-Animator [218] Video; Face PTA AnimateDiff Reference Feature Injection
@@
ConsisID [219] Video; Face PTA CogVideoX frequency decomposition
@@
MotionCharacter [220] Video; Face PTA VDM ID-Preserving Optimization; Motion Control Enhancement
@@
DreaMoving [126] Video; Face; Extra Conditions PTA SD Multi-feature Injection
Magic3D [138] 3D TTF Imagen; LDM Score Distillation Sampling
DreamBooth3D [140] 3D TTF Imagen Multi-stage Refinement
PAS [145] 3D PTA SD Text-to-3D-Pose
StyleAvatar3D [146] 3D PTA LDM Multi-view Alignment
AvatarBooth [147] 3D TTF SD Dual Model Fine-tuning
MVDream [139] 3D TTF SD 2.1 Score Distillation Sampling
Consist3D [141] 3D TTF SD Token Embedding Enhancement
Animate124 [143] 3D TTF SD 1.5 Multi-stage Refinement
Dream-in-4D [144] 3D TTF SD 2.1 Multi-stage Refinement
TextureDreamer [142] 3D TTF LDM Texture Extraction
@@
TIP-Editor [221] 3D TTF SD Multi-stage Refinement; Mask-assisted Generation
Anti-DreamBooth [148] Attack and Defense TTF SD 2.1 Perturbation Learning
@@
Concept Censorship [149] Attack and Defense TTF SD 1.4 Trigger Injection
Huang et al. [222] Attack and Defense TTF SD backdoor attack
Continual Diffusion [223] Others TTF SD Continual Learning
SVGCustomization [150] Others TTF SD 1.5 Fine-tuning and Alignment
@@
StitchDiffusion [151] Others TTF LDM Parameter-efficient Fine-tuning
MC-TI [224] Others TTF SD 1.5 Regularization

### v5

@@ 9 Conclusion
$\lambda$-eclipse [222patel2024lambda] Multiple Subjects PTA Kandinsky v2.2 Contrasstive Learning
MS-Diffusion [220personalizationms] Multiple Subjects PTA SD Attention-based Operation
ReVersion [63huang2023reversion] High-level Semantic TTF SD Regularization
Inv-ReVersion [159zhang2024inv] High-level Semantic TTF SD 1.5 Regularization
CusConcept [215xu2024cusconcept] High-level Semantic TTF SD 2.1 Regularization
@@
ADI [43huang2023learning] High-level Semantic TTF SD 2.1 Mask-assisted Generation
PhotoSwap [15gu2024photoswap] Extra Conditions TTF SD 2.1 Attention-based Operation
PE-VITON [38zhang2023two] Extra Conditions TTF SD Shape and Texture Control
@@
Layout-Control [152chen2024training] Extra Conditions TTF SD Attention-based Operation
SwapAnything [155gu2024swapanything] Extra Conditions TTF SD 2.1 Mask-assisted Generation
PE-VITON [38zhang2023two] Extra Conditions TTF SD Shape and Texture Control
@@
Viewpoint Control [163kumari2024customizing] Extra Conditions TTF SDXL 3D Feature Incorporation
@@
Prompt-Free Diffusion [110xu2023prompt] Extra Conditions PTA SD 2.0 Reference Feature Injection
Uni-ControlNet [111zhao2024uni] Extra Conditions PTA SD Multi-feature Injection
ViscoNet [93cheong2023visconet] Extra Conditions PTA SD Multi-feature Injection
@@
Context Diffusion [112najdenkoska2023context] Extra Conditions PTA LDM Multi-feature Injection
FreeControl [113mo2023freecontrol] Extra Conditions PTA SD 1.5; SD 2.1; SDXL Fusion Guidance
@@
Li et al. [168li2024tuning] Extra Conditions PTA SD Attention-based Operation; Mask-assisted Generation
@@
Tune-A-Video [150wu2023tune] Video TTF SD 1.4 Parameter-efficient Fine-tuning
Gen-1 [148esser2023structure] Video TTF LDM Multi-feature Injection
Make-A-Protagonist [135zhao2023make] Video TTF LDM Attention-based Operation
Animate-A-Story [136he2023animate] Video TTF SD Retrieval-augmented Paradigm
MotionDirector [101zhao2023motiondirector] Video TTF ZeroScope; ModelScope Disentanglement Approach
LAMP [102wu2023lamp] Video TTF SDXL; SD 1.4 First-frame Conditioned Pipeline
VMC [140jeong2023vmc] Video TTF Show-1 Parameter-efficient Fine-tuning
SAVE [103song2023save] Video TTF VDM Attention-based Operation
Customizing Motion [104materzynska2023customizing] Video TTF ZeroScope Regularization
DreamVideo [142wei2023dreamvideo] Video TTF ModelScope Disentanglement Approach
MotionCrafter [106zhang2023motioncrafter] Video TTF VDM Disentanglement Approach
Customize-A-Video [172ren2024customize] Video TTF ModelScope Multi-stage Refinement
Magic-Me [173ma2024magic] Video TTF SD 1.5; AnimateDiff Multi-stage Refinement
CustomTTT [202bi2024customttt] Video TTF VDM Fine-tuning and Alignment
PersonalVideo [208li2024personalvideo] Video; Face TTF AnimateDiff Identity Injection
@@
CustomVideo [145wang2024customvideo] Video; Multiple Subjects TTF ZeroScope Data Augmentation; Attention-based Operation
VideoDreamer [138chen2023videodreamer] Video; Multiple Subjects TTF SD 2.1 Data Augmentation
VideoAssembler [82zhao2023videoassembler] Video PTA VidRD Reference Feature Injection
VideoBooth [83jiang2023videobooth] Video PTA VDM Reference Feature Injection
VideoMaker [201wu2024videomaker] Video PTA VDM Reference Feature Injection
SUGAR [204zhou2024sugar] Video PTA VDM Attention-based Operation
StyleCrafter [212liu2024stylecrafter] Video; Style PTA SD 1.5 Reference Feature Injection
ID-Animator [158he2024id] Video; Face PTA AnimateDiff Reference Feature Injection
@@
ConsisID [207yuan2024identity] Video; Face PTA CogVideoX frequency decomposition
@@
MotionCharacter [206fang2024motioncharacter] Video; Face PTA VDM ID-Preserving Optimization; Motion Control Enhancement
@@
DreaMoving [105feng2023dreamoving] Video; Face; Extra Conditions PTA SD Multi-feature Injection
Magic3D [120lin2023magic3d] 3D TTF Imagen; LDM Score Distillation Sampling
DreamBooth3D [121raj2023dreambooth3d] 3D TTF Imagen Multi-stage Refinement
PAS [122azadi2023text] 3D PTA SD Text-to-3D-Pose
StyleAvatar3D [123zhang2023styleavatar3d] 3D PTA LDM Multi-view Alignment
AvatarBooth [124zeng2023avatarbooth] 3D TTF SD Dual Model Fine-tuning
MVDream [126shi2023mvdream] 3D TTF SD 2.1 Score Distillation Sampling
Consist3D [127ouyang2023chasing] 3D TTF SD Token Embedding Enhancement
Animate124 [128zhao2023animate124] 3D TTF SD 1.5 Multi-stage Refinement
Dream-in-4D [130zheng2023unified] 3D TTF SD 2.1 Multi-stage Refinement
TextureDreamer [133yeh2024texturedreamer] 3D TTF LDM Texture Extraction
@@
TIP-Editor [134zhuang2024tip] 3D TTF SD Multi-stage Refinement; Mask-assisted Generation
Anti-DreamBooth [13van2023anti] Attack and Defense TTF SD 2.1 Perturbation Learning
@@
Concept Censorship [50zhang2023backdooring] Attack and Defense TTF SD 1.4 Trigger Injection
Huang et al. [167huang2024personalization] Attack and Defense TTF SD backdoor attack
Continual Diffusion [64smith2023continual] Others TTF SD Continual Learning
SVGCustomization [36zhang2023text] Others TTF SD 1.5 Fine-tuning and Alignment
@@
StitchDiffusion [25wang2024customizing] Others TTF LDM Parameter-efficient Fine-tuning
MC-TI [214wang2024multi] Others TTF SD 1.5 Regularization
@@ Acknowledgments
Xulu Zhang received the B.E., and M.S. degrees from Sichuan University in 2019 and 2022. He is currently working towards the Ph.D degree from the Department of Computing at the Hong Kong Polytechnic University. His research interests include image generation and active learning.
E-mail: compxulu.zhang@connect.polyu.hk
ORCID iD: 0000-0003-2473-460X
Xiaoyong Wei received the Ph.D. degree in computer science from City University of Hong Kong, China in 2009, and has worked as a postdoctoral fellow in the University of California, Berkeley, USA from December, 2013 to December, 2015. He has been a professor and the head of Department of Computer Science, Sichuan University, China since 2010. He is an adjunct professor of Peng Cheng Laboratory, China, and a visiting professor of Department of Computing, Hong Kong Polytechnic University. He is a senior member of IEEE, and has served as an associate editor of Interdisciplinary Sciences: Computational Life Sciences since 2020, the program Chair of ICMR 2019, ICIMCS 2012, and the technical committee member of over 20 conferences such as ICCV, CVPR, SIGKDD, ACM MM, ICME, and ICIP.
His research interests include multimedia computing, health computing, machine learning and large-scale data mining.
E-mail: x1wei@polyu.edu.hk (Corresponding Author)
ORCID iD: 0000-0002-5706-5177
Wentao Hu received the B.E., and M.S. degrees from Shandong University and Sun Yat-sen University, in 2021 and 2024. He is currently working towards the Ph.D degree from the Department of Computing at the Hong Kong Polytechnic University. His research interests include image generation and 3D reconstruction.
E-mail: wayne-wt.hu@connect.polyu.hk
ORCID iD: 0000-0002-2071-9341
Jinlin Wu is an Assistant Research Fellow at the Institute of Automation, Chinese Academy of Sciences (CAS). He received his Ph.D. from the University of Chinese Academy of Sciences in 2022 and his bachelor’s degree from the University of Electronic Science and Technology of China in 2017. His research interests include object detection, image recognition, and video understanding. He has served as the principal investigator of a National Natural Science Foundation of China (NSFC) Youth Science Fund project and has participated in several other NSFC-funded projects. Dr. Wu has accumulated a solid research foundation in areas such as video analysis for security and medical video understanding. He has published over 30 high-quality academic papers, with more than 500 citations.
E-mail: jinlin.wu@cair-cas.org.hk
ORCID iD: 0000-0001-7877-5728
Jiaxin Wu is a Postdoctoral Fellow in the Department of Computing at The Hong Kong Polytechnic University. She earned her Ph.D. in Computer Science from the City University of Hong Kong. Her research interests include multimedia retrieval, AI for Science (AI4Science), and natural language processing (NLP).
E-mail: jiaxwu@polyu.edu.hk
ORCID iD: 0000-0003-4074-3442
Wengyu Zhang is an Undergraduate Student in the Department of Computing at The Hong Kong Polytechnic University. His research interests include natural language processing (NLP), AI for Science (AI4Science), and Graph Learning.
E-mail: wengyu.zhang@connect.polyu.hk
ORCID iD: 0009-0001-2347-4183
Zhaoxiang Zhang received his BSc degree in Department of Electronic Science and Technology from University of Science and Technology of China, and the PhD degree in Pattern Recognition and Intelligent Systems from the Institute of Automation, Chinese Academy of Sciences in 2004 and 2009, respectively. From 2009 to 2015, he worked as a Lecturer, Associate Professor, and later the deputy director of Department of Computer Application Technology at the Beihang University. Since July 2015, Dr. Zhang has joined the Institute of Automation, Chinese Academy of Sciences, where he is currently a professor. His major research interests include pattern recognition, computer vision and machine learning. Recently, he specifically focuses on deep learning models, biologically-inspired visual computing and human-like learning, and their applications on human analysis and scene understanding. He has published more than 200 papers in international journals and conferences, including reputable international journals such as IEEE TPAMI, IJCV, JMLR, IEEE TIP, IEEE TCSVT, and top-level international conferences like CVPR, ICCV, NIPS, ECCV, ICLR, AAAI, IJCAI and ACM MM. He has won the best paper awards in several conferences and championships in international competitions and his research has won the ‘Technical Innovation Award of the Chinese Association of Artificial Intelligence’. He has served as the PC Chair or area chair of many international conferences like CVPR, ICCV, AAAI, IJCAI, ACM MM, ICPR and BICS. He is serving or has served as associate editor of reputable international journals like IJCV, IEEE T-CSVT, IEEE T-BIOM, Pattern Recognition and NeuroComputing.
E-mail: zhaoxiang.zhang@ia.ac.cn
ORCID iD: 0000-0003-2648-3875
Zhen Lei received the B.S. degree in automation from the University of Science and Technology of China, in 2005, and the Ph.D. degree from the Institute of Automation, Chinese Academy of Sciences, in 2010, where he is currently a professor. He is IEEE Fellow, IAPR Fellow and AAIA Fellow. He has published over 200 papers in international journals and conferences with 33000+ citations in Google Scholar and h-index 86. He was the program co-chair of IJCB2023, was competition co-chair of IJCB2022 and has served as area chairs for several conferences and is associate editor for IEEE Trans. on Information Forensics and Security, IEEE Transactions on Biometrics, Behavior, and Identity Science, Pattern Recognition, Neurocomputing and IET Computer Vision journals. His research interests are in computer vision, pattern recognition, image processing, and face recognition in particular. He is the winner of 2019 IAPR Young Biometrics Investigator Award.
E-mail: zhen.lei@ia.ac.cn (Corresponding Author)
ORCID iD: 0000-0002-0791-189X
Qing Li is currently a Chair Professor (Data Science) and the Head of the Department of Computing, the Hong Kong Polytechnic University. Formerly, he was the founding Director of the Multimedia software Engineering Research Centre (MERC), and a Professor at City University of Hong Kong where he worked in the Department of Computer Science from 1998 to 2018. Prior to these, he has also taught at the Hong Kong University of Science and Technology and the Australian National University (Canberra, Australia). Prof. Li served as a consultant to Microsoft Research Asia (Beijing, China), Motorola Global Computing and Telecommunications Division (Tianjin Regional Operations Center), and the Division of Information Technology, Commonwealth Scientific and Industrial Research Organization (CSIRO) in Australia. He has been an Adjunct Professor of the University of Science and Technology of China (USTC) and the Wuhan University, and a Guest Professor of the Hunan University (Changsha, China) where he got his BEng. degree from the Department of Computer Science in 1982. He is also a Guest Professor (Software Technology) of the Zhejiang University (Hangzhou, China) – the leading university of the Zhejiang province where he was born.
Prof. Li has been actively involved in the research community by serving as an associate editor and reviewer for technical journals, and as an organizer/co-organizer of numerous international conferences. Some recent conferences in which he is playing or has played major roles include APWeb-WAIM’18, ICDM 2018, WISE2017, ICDSC2016, DASFAA2015, U-Media2014, ER2013, RecSys2013, NDBC2012, ICMR2012, CoopIS2011, WAIM2010, DASFAA2010, APWeb-WAIM’09, ER’08, WISE’07, ICWL’06, HSI’05, WAIM’04, IDEAS’03,VLDB’02, PAKDD’01, IFIP 2.6 Working Conference on Database Semantics (DS-9), IDS’00, and WISE’00. In addition, he served as a programme committee member for over fifty international conferences (including VLDB, ICDE, WWW, DASFAA, ER, CIKM, CAiSE, CoopIS, and FODO). He is currently a Fellow of IEEE and IET/IEE, a member of ACM-SIGMOD and IEEE Technical Committee on Data Engineering. He is the chairperson of the Hong Kong Web Society, and also served/is serving as an executive committee (EXCO) member of IEEE-Hong Kong Computer Chapter and ACM Hong Kong Chapter. In addition, he serves as a councilor of the Database Society of Chinese Computer Federation (CCF), a member of the Big Data Expert Committee of CCF, and is a Steering Committee member of DASFAA, ER, ICWL, UMEDIA, and WISE Society.
E-mail: qing-prof.li@polyu.edu.hk
ORCID iD: 0000-0003-3370-471X